# ELASTIC: Efficient Once For All Iterative Search for Object Detection on Microcontrollers

### v2

@@ IV-D Iterative Search Space Refinement
To evaluate how ELASTIC improves the underlying search space over time, we conduct a series of random sampling experiments across both SVHN and PascalVOC. Specifically, we compare the distribution of model accuracies under three search conditions: (1) the original joint search space, (2) the progressive head-only search space, and (3) the refined head space produced by ELASTIC after iterative search.
Figure 4 shows the accuracy scatter plots and means of 100 randomly sampled architectures from each space. On SVHN, the joint space produces models with a mean accuracy of 44.61%, while progressive head-only search improves this to 67.38%. ELASTIC further enhances the head space, achieving a mean accuracy of 72.25% — a 4.87% increase over progressive search. A similar trend is observed on PascalVOC, where the mean accuracy improves from 4.99% (joint) to 27.89% (progressive), and then to 28.32% with ELASTIC.
To further examine the progression of search space quality, we track mean and variance over time using random samples taken at key iterations of ELASTIC. As shown in Figure 5 and Table III, the mean accuracy on SVHN improves from 67.38% at iteration 0 to 70.68% at iteration 5, while variance decreases, indicating a more stable and optimized search space. Similar patterns are observed on PascalVOC, where the mean increases from 27.89% to 28.32%, with variance decreasing from to .
IV-E Deployment on MAX78000/MAX78002 Platform
The MAX78000/MAX78002 platform integrates a dedicated CNN accelerator but imposes strict architectural and memory constraints. For example, the MAX78000 includes a 432 KB weight memory limit, 32 KB of data memory, support restricted to and convolution kernels, a maximum of 32 layers, and up to 1024 channels per layer while the MAX78002 includes a 2.3 MB weight memory limit, 80 KB of data memory, and a maximum of 128 layers.
Despite these limitations, our ELASTIC-derived models demonstrate substantial improvements over Analog Devices’ baselines [21]. On PascalVOC, ELASTIC reduces energy by 71.6%, lowers latency by 2.40, and improves mAP by +6.99, while remaining within the on-chip memory budget. On VGGFace2 and SVHN, ELASTIC achieves consistent efficiency gains, reducing energy by 20.1% and 40.5%, power by 30.1% and 42.8%, and latency by 5.1%and 7.1%, respectively, with mAP gains of +2.38% and +4.50%.
Figure 6 compares ELASTIC-derived architectures with manually scaled TinierSSD [21] baselines in terms of accuracy and FLOPs. ELASTIC consistently achieves higher accuracy while significantly reducing FLOPs—up to 66%— at similarly sized models, highlighting the benefits of hardware-aware NAS.
@@ V Conclusion
Through extensive experiments on multiple datasets and deployment on real microcontroller platforms, we demonstrated that ELASTIC discovers architectures that are both more accurate and more efficient than manually designed and standard NAS baselines. These results highlight the effectiveness of iterative, hardware-aware NAS in advancing the accuracy–efficiency frontier for resource-constrained edge devices.

### v3

@@ IV-D Iterative Search Space Refinement
To evaluate whether ELASTIC improves the effective search space, we randomly sample 100 architectures from three matched spaces: (1) joint backbone–neck/head, (2) progressive head-only (fixed backbone), and (3) ELASTIC-refined head-only (after re-optimization) in Figure 4. On SVHN, the mean mAP increases from 44.61% (joint) to 67.38% (progressive) and to 72.25% with ELASTIC (+4.87% vs. progressive). On PascalVOC, the mean improves from 4.99% (joint) to 27.89% (progressive) and to 28.32% with ELASTIC.
We further track refinement over iterations by repeating the sampling at iterations 0, 3, 5 (Fig. 5, Table III), treating iteration 0 as the progressive baseline. On SVHN, the mean mAP rises from 67.38% (iter. 0) to 70.68% (iter. 5), with variance changing from 0.136 to 0.199. On PascalVOC, the mean increases from 27.89% to 28.32%, while variance decreases from to . Overall, iterative revisiting yields a stronger induced space than progressive partitioning and improves stability on PascalVOC.
IV-E Deployment on MCU Platforms
The MAX78000/MAX78002 MCUs integrate a CNN accelerator but impose tight constraints on kernel sizes, layer counts, and memory. Despite these limits, ELASTIC-derived models improve the efficiency–accuracy trade-off over TinierSSD/FPN baselines [9] while remaining within on-chip budgets.
We highlight these results in Table IV. On PascalVOC, energy decreases by 71.6%, latency by 2.40, power by 36.1%, and mAP increases by +6.99 on MAX78002. On TrashNet, ELASTIC follows the same pattern on MAX78002: energy decreases by 71.6%, latency by 2.40, power by 36.1%, and mAP increases by +10.2. ELASTIC also generalizes across tasks: on VGGFace2, it reduces energy and power by 20.1% and 30.1%, respectively, while improving mAP by +2.38; on SVHN, it reduces energy and power by 40.5% and 42.8%, respectively, with a +4.50 mAP gain and a slight latency improvement. We further evaluate cross-device portability on STM32F746. On PascalVOC, the scaled ELASTIC model reduces latency slightly from 494.8 ms to 490.6 ms while improving mAP from 9.12 to 19.0. On TrashNet, the scaled ELASTIC model reduces latency slightly from 494.8 ms to 490.6 ms while improving mAP from 54.03 to 73.9. On VGGFace2, ELASTIC reduces latency by 27.8% and improves mAP by +1.9. On SVHN, ELASTIC improves mAP by +4.4.
Since the two scaled STM32F746 models differ in their maximum deployable input resolution, we also compare them at both and in Table V. The scaled ai87-fpndetector baseline can be deployed only up to , while the scaled ELASTIC model remains deployable at ; at , the baseline encounters out-of-memory due to activation SRAM limits on STM32F746. Under matched resolution, ELASTIC still outperforms the baseline on both datasets, improving mAP from 9.12 to 19.0 on PascalVOC and from 54.03 to 73.9 on TrashNet, while maintaining nearly identical latency (490.6 ms vs. 494.8 ms). Moreover, ELASTIC supports the higher resolution and achieves 40.1 mAP on PascalVOC and 86.0 mAP on TrashNet, further highlighting that the searched architecture is not only more accurate but also more deployable under tight STM32F746 memory constraints.
Figure 6 compares ELASTIC architectures to scaled TinierSSD baselines [9] and shows ELASTIC attains higher accuracy at lower compute (66% FLOPs reduction), underscoring the value of hardware-aware NAS for MCU deployment.
| Model | Dataset | Device | Params | Resolution | Latency (ms) | mAP (%) |
| ai87-fpndetector [9] | PascalVOC | STM32746 | 0.82M | 128128 | N/A | OOM |
| ai87-fpndetector [9] | PascalVOC | STM32746 | 0.82M | 100100 | 494.8 | 9.12 |
| ELASTIC (OURS) | PascalVOC | STM32746 | 0.89M | 128128 | 825.55 | 40.1 |
| ELASTIC (OURS) | PascalVOC | STM32746 | 0.89M | 100100 | 490.6 | 19.0 |
| ai87-fpndetector [9] | TrashNet | STM32746 | 0.82M | 128128 | N/A | OOM |
| ai87-fpndetector [9] | TrashNet | STM32746 | 0.82M | 100100 | 494.8 | 54.03 |
| ELASTIC (OURS) | TrashNet | STM32746 | 0.89M | 128128 | 825.55 | 86.0 |
| ELASTIC (OURS) | TrashNet | STM32746 | 0.89M | 100100 | 490.6 | 73.9 |
@@ V Conclusion
Through extensive experiments on multiple datasets and deployment on real microcontroller platforms, we demonstrated that ELASTIC discovers architectures that are more accurate and more efficient than manually designed and standard NAS baselines. These results highlight the effectiveness of iterative, hardware-aware NAS in advancing the accuracy–efficiency frontier for resource-constrained edge devices.